You are working in a biology lab during your summer break. Your supervisor asks you to perform an experiment to find the effective spring constant (in N/m) of a partial molecule of DNA (deoxyribonucleic acid). You perform experiments and find that a single straight portion of a DNA molecule is 2.29 m long. You then perform an activity that charges the ends of the molecule; each end becomes singly ionized: negative on one end, positive on the other. After the ends are charged, the molecule shrinks by 1.16% of its length. N/m Noed Help?
Answer #1

We know that

F = kx => k = F/x

x = 1.16 % of 2.29 um = 0.027 um

F = k q1q2/r^2

q1 = q2 = e = 1.6 x 10^-19

F = 9 x 10^9 x (1.6 x 10^-19)/(2.29 x 10^-6)^2 = 274.6 N

k = 274.6/(0.027 x 10^-6) = 1.02 x 10^10 N/m

Hence, k = 1.02 x 10^10 N/m